• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Snyder

    Log entry time 08:12:05 on July17,2004

    Entry number 128749

    keyword=Owl Shift Summary

    00:40 Moller measurement completed by Eugene and his crew
    00:43 Run 2906 started; production
    01:39 Run 2906 stopped; IHWP switched to OUT; bleedthrough measurement done (0.21% bleedthrough)
    01:43 Run 2907 started; 50k production events

    02:15 noticed ROC 25 & ROC 26 were not responding
    tried resetting coda, but both ROCS did not boot up
    paged Bryan Moffit to come in and reset network 192 switch

    05:00 returned to beam permit and started run 2908;production
    --MCC has IOC issue so most of run is pedestal
    06:20 run 2908 stopped
    06:21 run 2909 started; pedestal--> 30k events
    6:39 run 2910 started; production
    07:39 run 2910 stopped
    07:39 run 2911 started